Consider the following reaction that was run at STP, if 200 mL of nitrogen are reacted with 150 mL of oxygen what is the maximum number of grams of NO2 that could be produced?
N2+2O2–>2NO2
1
Expert's answer
2020-11-09T14:03:52-0500

N2 + 2 O2 = 2 NO2

Quantity of N2: 0.200 L / (22.4 L/mol) = 0.0089 mol

Quantity of O2: 0.150 L / (22.4 L/mol) = 0.0067 mol

According to the reaction, nitrogen is in excess.

The quantity of NO2 is equal to the quantity of oxygen: 0.0067 mol

The mass of NO2: 0.0067 mol * (46 g/mol) = 0.3 g

Answer: 0.3 g